OpenStudy (anonymous):

solve for x. 3y - 6/x = 12

OpenStudy (tkhunny):

Have you forgotten your Order of Operations? You have if you intended to write \(\dfrac{3y-6}{x} = 12\)

OpenStudy (anonymous):

i know my Order of Operation but i dont know how to solve this problem

OpenStudy (tkhunny):

You did not answer my question. Did you intend to write \(3y - \dfrac{6}{x} = 12\) or not? That is what you wrote.

OpenStudy (anonymous):

yes thats what i intended to write

OpenStudy (tkhunny):

Perfect. You would be surprised how many don't get it at all! It makes me a little cautious. Good work being one of the few!! #1 thing to do is consider the Domain. \(x \ne 0\). Do you see why this is so?

OpenStudy (anonymous):

why is it \[x \neq 0\]

OpenStudy (tkhunny):

With that x in the denominator, that term would blow up. Writing it down FIRST may save us headaches later. We may manipulate the equation in a way where this information is lost. It pays big dividends to write it down first. #2 thing to do is look for ways to simplify your life. In this case, it appears 3's are wild. Let's divide each term by 3.

OpenStudy (anonymous):

meaning?

OpenStudy (tkhunny):

Meaning what? Divide everything by 3. This will be a helpful simplification. Now we have : \(y - \dfrac{2}{x} = 4\)

OpenStudy (anonymous):

whats the answer?

OpenStudy (tkhunny):

It doesn't work that way here. We need to see YOUR work. I gave you the first two hints.

OpenStudy (anonymous):

so do multiply y and x?

OpenStudy (tkhunny):

It may be tempting to multiply by x, since that may get it out of the denominator, but that will not help us. It will create a different problem. Persanally, I would add 2/x to both sides. Let's see what you get.

OpenStudy (anonymous):

i dont know how to do that.. i need help

OpenStudy (tkhunny):

\(y = 4 + \dfrac{2}{x}\) Next, I would subtract 4 from both sides. You have to show me this one.

OpenStudy (anonymous):

-4y=2/x ?

OpenStudy (tkhunny):

Needs a little work. 4 is now multiplied by y. It should be subtraction. Give it another go.